What is the hydronium-ion concentration of a 0.03 F HCN solution?


1
Expert's answer
2021-03-18T02:51:24-0400

According to the equation:

Ka = [H+][CN-] / [HCN]

From here:

[H+][CN-] = Ka × [HCN]

[H+]² = Ka × [HCN]

[H+] = (Ka × [HCN])1/2 = 4.5 × 10-10 × 0.03 F = 3.7 × 10-4 F


Answer: 3.7 × 10-4 F